Association of attributes :

The patients reactions to the treatment are recorded in the following table. Test the hypothesis that the drug is no better than sugar pills foru cring colds. Let a =0.05.


Category


Drug, Sugar


Helped

52, 44

Harmed

10, 12

No Effect

20, 26


1
Expert's answer
2021-01-17T14:54:55-0500

The expected frequencies are


"\frac{row\ total*column\ total}{grand\ total}"


"Drug.helped= \frac{(52+44)(52+10+20)}{52+10+20+44+12+26}= 48"


"Drug.harmed= \frac{(10+12)(52+10+20)}{52+10+20+44+12+26}= 11"


"Drug.no.effect= \frac{(20+26)(52+10+20)}{52+10+20+44+12+26}= 23"


"Sugar.helped= \frac{(52+44)(44+12+26)}{52+10+20+44+12+26}= 48"


"Sugar.harmed= \frac{(10+12)(44+12+26)}{52+10+20+44+12+26}= 11"


"Sugar.no.effect= \frac{(20+26)(44+12+26)}{52+10+20+44+12+26}= 23"

The degrees of freedom for the chi square coefficient is


"(3-1)(2-1)=2"

The chi square test statistic is given by


"X^2=\sum\frac{(observed-expected)^2}{expected}"

"X^2=\frac{(52-48)^2}{48}+\frac{(10-11)^2}{11}+\frac{(20-23)^2}{23}+ 
\frac{(44-48)^2}{48}+\frac{(12-11)^2}{11}+\frac{(26-23)^2}{23}"

"X^2=1.6311"

The p value corresponding to the chi square score of "1.6311" and a level of significance of"0.05" is "0.4424" . We therefore conclude that there is no significant evidence to conclude that there is a difference in sugar and the drug.